JuniperBoi:

A line has a slope of 1 and passes through the point (–8, 0) . What is its equation in slope -intercept form?

justus:

use the point slope formula

justus:

y – y1 = m(x – x1)

JuniperBoi:

Ok, thanks. I'll try it out, this is just difficult T^T lol

justus:

Sure. Nah it's very simple. You can do it!

JuniperBoi:

c:

justus:

y+0=1(x+8)